Find the future value of $1125 at 5.2% for four years if the account pays simple innterest
Answer by Aldorozos(172)

FV = PV(1+r)^n
PV = 1125
r = 5.2%/12 (monthly
n = 4*12 = 12
Here we are assuming monthly interest. If it is only annual interest then we have to set r = 5.2% and n = 4
You may use the following to calculate the future value:
http://www.investopedia.com/calculator/fvcal.aspx
Here is for annual
$1,377.89